# Smooth Deep Saliency

### v1

@@ 4.2.1 Accuracy and Prediction Difference
4.2.2 Total Variation
Table 4 shows the total variation (TV) of the saliency map for both the original and surrogate models, both for DeepLift and Gradient. We report the mean and standard deviation over the validation data. The results have some stochasticity since we add gaussian noise to the input (we use SmoothGrad where we run the saliency method 20 times each time adding gaussian noise with a standard deviation of 0.2 to the input, and return the mean of those 20 runs as the saliency map). Layer1, Layer2 and Layer3 denote the last layer at that resolution stage. This means Layer1 is the last layer where the input has been downsampled twice, Layer2 is the last layer where it has been downsampled three times, and Layer3 is the last layer where it has been downsampled four times. We use two saliency methods: DeepLift and Gradient (which is simply the gradient with respect to the hidden layer).
With our bilinear surrogate models, we observe a reduction of TV of the saliency map compared to the original model, through all models and layers. For DeepLift we observe an average reduction in TV of 21.6%, with a minimum of 9.6% in Layer3 of ResNet50 trained on ImageNet1K, and a maximum of 47.8% in Layer3 of ResNet34 trained on Digipath. Even greater reductions can be observed with the Gradient method. There we observe an average reduction in TV of 31.5%, with a minimum of 18.2% in Layer3 of ResNet50 trained on ImageNet1K, and a maximum of 58.0% in Layer3 of ResNet34 trained on Digipath.
Figure 4 shows the total variation for the original and bilinear surrogate model for the ResNet34, ResNet50 and ResNet50 Robust models on ImageNet1K when computing the saliency map using the Gradient method. TV is shown on the y-axis, whereas the layer is on the x-axis. The figure shows the total variation for all the layer blocks starting after the max pool (at two downsamplings away from the input) till the layer before the lowest resolution stage. We do not evaluate earlier layers since we want to compute saliency towards deeper layers. And we do not evaluate layers at the lowest resolution stage since there is no convolutional downsampling coming after them, and therefore no checkerboard noise in the gradient. Layer1, Layer2 and Layer3 from Table 4 correspond to $1\_2$, $2\_3$ respectively $3\_5$ in Figure 4. We observe a reduction in TV for our bilinear surrogate models compared to the original models through all layers. The backward hook method has slightly lower TV compared to the bilinear surrogate method, and the forward hook method has the lowest TV.
Figure 5: Saliency map from DeepLift from original model (first 4 images from the left), bilinear surrogate model (second 4 images from the left), backward hooked model (third 4 images) and forward hooked model (last 4 images). For the robust ResNet50. From top to bottom shows the original input image, saliency map for the input layer, Layer1, Layer2, Layer3.
@@ 4.2.2 Total Variation
Figure 8: Saliency map from DeepLift on Digipath data, zoomed in on Figure 7. From left to right is: Input image, saliency map for the original model in the input layer, saliency map of the surrogate model for Layer1, saliency map of the original model for Layer1.
4.2.3 Case Studies
@@
An example for the usefulness of using saliency maps in hidden layers, is shown in Figure 7 and Figure 8. Figure 7 shows the saliency map for a single input image, for both the original model (bottom two rows) and the bilinear surrogate model (top two rows) for the input layer, Layer1, Layer2, Layer3, Layer4 (from left to right). The top of the two rows is the saliency map, the bottom is overlaying it with the input image. The saliency map for the input layer is highlighting edges and noise, and does not give much information about which structures the network uses to make its prediction. Similarly, the saliency map for Layer4, which is the last convolutional layer (in the encoder of the model), is also not very helpful here, because it is just highlighting the whole tumor. Since the model is a semantic segmentation network, this does not add more information than the prediction already gives.
@@
Figure 8 only shows the center 512x512 pixels of the 1536x1536 images of Figure 7 for the input layer and Layer1. There we can see the highlighted cell nuclei more clearly.
Figure 5 shows saliency maps for the robust ResNet50. From left to right in pairs of 4 images it shows the saliency maps for the original model, the bilinear surrogate model, the backward hooked model and the forward hooked model. From top to bottom it shows the original input, saliency map for the input, Layer1, Layer2 and Layer3. The saliency map at the input is smooth in all cases. But for the deeper layers there is a strong checkerboard noise pattern for the original model, whereas the maps for the other methods (bilinear surrogate, backward hooked model, forward hooked model) are smooth.
Something similar can be observed for Figure 6. The figure shows saliency maps for the ResNet18 Camelyon16 for the original (left 4 images) and bilinear surrogate model (right 4 images) for Camelyon16 samples. From top to bottom it shows the input image, saliency map for the input, for Layer1, Layer2, Layer3. The model classified all 4 images as tumor. The input layer is noisy and difficult to interpret. In the deeper layers, the original model exhibits checkerboard noise in its saliency maps, whereas the saliency maps for the bilinear surrogate model are smoother and easier to interpret. In Layer1 we can see cell nuclei being outlined, whereas deeper in Layer2 we can see the segmentations of whole cell nuclei.
4.3 Additional Discussion
While the forward hook method has the lowest TV, and therefore smoothest saliency maps, it has considerably lower accuracy compared to the original model and its predictions differ rather strongly. The bilinear surrogate method closely matches the accuracy of the original model, while considerably reducing the TV and the checkerboard noise. The backward hook method also considerably reduces the TV and needs no training of a surrogate model, but here it is unclear how faithful the saliency map is to the original model.
@@ 5 Limitations
6 Conclusion and Future Work
We successfully investigated three methods to remove checkerboard noise of saliency maps in hidden layers of models that utilize convolutional downsampling. The bilinear surrogate method closely matches the accuracy and predictions of the original model, while reducing the noise in the saliency map, measured by TV. We observed a reduction in TV through all measured layers. While the forward hook method also removes the checkerboard noise, it has a lower accuracy and its predictions differ more strongly from the original model. Overall we recommend first trying the backward hook method, which removes the checkerboard noise and requires no additional training, then if the saliency map from hidden layers gives useful results, train a bilinear surrogate model.
For future work, it would be interesting to check whether we can use the explanations gained from the bilinear surrogate model to explain the original model, but more experiments specifically to validate this would be needed.
7 Impact Statement
This paper presents work whose goal is to improve transparency of machine learning models. Apart from the many potential societal consequences of generally advancing the field of machine learning, improving transparency of the models can help with their acceptance in a broad sense.
@@ 8 Acknowledgements
R.H. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - project number 459360854. M. S. and D. O. B. acknowledge the financial support by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) and the European Social Fund (ESF) within the EXIST transfer of research project ”aisencia” and the financial support by the Federal Ministry of Education and Research (BMBF) within the T!Raum project ”MOIN - MUKIDerm”.

### v3

@@ 4.4 Randomize Model Weights
Table 5: Percentage reduction in total variation of the saliency map compared to the original model.
Method Network S B F
Grad ResNet34 11.1 18.7 21.5
ResNet50 8.2 13.2 16.3
ResNet50 Robust 14.1 18.7 20.7
ResNet18 Camelyon16 16.9 27.2 28.7
Total 12.6 19.5 21.8
DeepLift Total 10.5 13.9 17.2
IG Total 8.6 12.2 15.7
4.5 Total Variation
In this section we report our results on the total variation experiments. Fig. 6 shows the total variation of the gradient throughout hidden layers of the ImageNet1K models (going deeper into the model from left to right). It shows the result for the original model, and the three methods used to reduce the checkerboard artifacts (bilinear surrogate, backward hook, forward hook). We can observe that all three methods reduce the total variation at every layer compared to using the original model.
Tab. 5 shows the percentage reduction in total variation of the saliency map compared to the original model. On the right side of the table S stands for bilinear surrogate, B for backward hook and F for forward hook. For the gradient we report both the total variation per model (averaged over all layers) as well as the total (averaged over all models). Through all models and methods we see the largest reduction in TV for the ResNet18 Camelyon16 model, and the smallest reduction in the ResNet50 model.
Figure 7: Saliency map from DeepLift from original model (first 4 images from the left), bilinear surrogate model (second 4 images from the left), backward hooked model (third 4 images) and forward hooked model (last 4 images). For the ResNet34. From top to bottom shows the original input image, saliency map for the input layer, Layer1, Layer2, Layer3.
@@ 4.5 Total Variation
Figure 10: Saliency map from DeepLift on Digipath data, zoomed in on Fig. 9. From left to right is: Input image, saliency map for the original model in the input layer, saliency map of the surrogate model for Layer1, saliency map of the original model for Layer1.
4.6 Case Studies
@@
An example for the usefulness of using saliency maps in hidden layers, is shown in Fig. 9 and Fig. 10. Fig. 9 shows the saliency map for a single input image, for both the original model (bottom two rows) and the bilinear surrogate model (top two rows) for the input layer, Layer1, Layer2, Layer3, Layer4 (from left to right). The top of the two rows is the saliency map, the bottom is overlaying it with the input image. The saliency map for the input layer is highlighting edges and noise, and does not give much information about which structures the network uses to make its prediction. Similarly, the saliency map for Layer4, which is the last convolutional layer (in the encoder of the model), is also not very helpful here, because it is just highlighting the whole tumor. Since the model is a semantic segmentation network, this does not add more information than the prediction already gives.
@@
Fig. 10 only shows the center 512x512 pixels of the 1536x1536 images of Fig. 9 for the input layer and Layer1. There we can see the highlighted cell nuclei more clearly.
Fig. 7 shows saliency maps for the ResNet34. From left to right in pairs of 4 images it shows the saliency maps for the original model, the bilinear surrogate model, the backward hooked model and the forward hooked model. From top to bottom it shows the original input, saliency map for the input, Layer1, Layer2 and Layer3. The saliency map at the input is smooth in all cases. But for the deeper layers there is a strong checkerboard noise pattern for the original model, whereas the maps for the other methods (bilinear surrogate, backward hooked model, forward hooked model) are smooth.
Something similar can be observed for Fig. 8. The figure shows saliency maps for the ResNet18 Camelyon16 for the original (left 4 images) and bilinear surrogate model (right 4 images) for Camelyon16 samples. From top to bottom it shows the input image, saliency map for the input, for Layer1, Layer2, Layer3. The model classified all 4 images as tumor. The input layer is noisy and difficult to interpret. In the deeper layers, the original model exhibits checkerboard noise in its saliency maps, whereas the saliency maps for the bilinear surrogate model are smoother and easier to interpret. In Layer1 we can see cell nuclei being outlined, whereas deeper in Layer2 we can see the segmentations of whole cell nuclei.
@@ 5 Limitations
6 Conclusion
We successfully investigated three methods to remove checkerboard noise of saliency maps in hidden layers of models that utilize convolutional downsampling. The bilinear surrogate method closely matches the accuracy and predictions of the original model, while reducing the noise in the saliency map, measured by TV. We observed a reduction in TV through all measured layers, with an average reduction of 10.5% for DeepLift. It performs slightly better under insertion metrics (2.14%) compared to the original model, but a bit worse under deletion metrics (11.4%). The backward hook method also considerably reduces the TV and needs no training of a surrogate model, but it performs slightly worse under insertion and deletion metrics compared to the bilinear surrogate. While the forward hook method has the lowest TV, and therefore smoothest saliency maps, it has considerably lower accuracy compared to the original model, its predictions differ rather strongly and it is not very faithful under insertion and deletion metrics. Overall we recommend first trying the backward hook method, which removes the checkerboard noise and requires no additional training. Then if the saliency map from hidden layers gives useful results, train a bilinear surrogate model.
@@ 7 Acknowledgements
R.H. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projectnumber 459360854 (DFG FOR 5347 Lifespan AI). M.S. and D.O.B. acknowledge the financial support by the German Federal Ministry of Economic Affairs and Climate Action (BMWK) and the European Social Fund (ESF) within the EXIST transfer of research project "aisencia" and the financial support by the Federal Ministry of Education and Research (BMBF) within the T!Raum-Initiative "#MOIN! Modellregion Industriemathematik" in the sub-project "MUKIDerm".